# Multivariate Causal Effects: a Bayesian Causal Regression Factor Model

### v2

@@ 3. Bayesian Causal Factor Model
where $\Phi(\cdot)$ is the Probit function and $\{a_{r}^{(tj)}({\bf x}_{i})\}_{r\geq 1}$ have Gaussian distributions whose mean is a linear combination of the $p$ covariates ${\bf x}_{i}$.
@@
This prior distribution for treatment-specific factors (3) allows us to introduce a latent categorical variable $S_{itj}$, for each factor $j\in\{1,\dots,J_{t}\}$ treatment $t\in\{0,1\}$ unit $i\in\{1,\dots.n\}$, describing clusters of units defined by heterogeneous responses to the treatment level $t$. Assuming $\Pr(S_{itj}=r)=\pi_{r}^{(tj)}({\bf x}_{i})$, then the prior distribution of the treatment-specific factors (3) can be rewritten, conditional on $S_{itj}$, as
(6) $\{{\bf l}_{itj}\mid S_{itj}=r,\boldsymbol{\omega}\}\sim\mathcal{N}(\eta_{r}^{(tj)},1/\tau_{r}^{(tj)}).$
Statistical Guarantees
@@
The proposed prior for treatment-specific factor scores offers several strengths, as outlined in Section The Role of the DDP in the Factor Scores. Additionally, it serves as a powerful tool for inferring information about the relationships among the multiple outcomes, as governed by unobserved features $U$. Specifically, the flexibility of the DDP distribution enables it to capture this unmeasured information while effectively controlling bias.
In particular, we consider the case where $U\sim\mathcal{N}(\mu_{u},\sigma_{u}^{2})$ with non-negative correlation $\sigma^{(lu)}$ between $U_{i}$ and $l_{itj}$, allowing us to derive closed-form expressions that clarify the relationships among the quantities involved.
(11) $\left[\begin{array}[]{c}{\bf l}_{itj}\mid{\bf x}_{i}\\ U_{i}\end{array}\right]\sim\sum_{r\geq 1}\pi_{r}^{(tj)}({\bf x}_{i})\mathcal{N}\left(\left[\begin{array}[]{c}\eta_{r}^{(tj)}\\ \mu_{u}\end{array}\right],\Sigma_{r}^{(LU)}\right),$
(14) $\Sigma_{r}^{(LU)}=\left[\begin{array}[]{cc}1/\tau_{r}^{(tj)}&\sigma^{(lu)}\\ \sigma^{(lu)}&\sigma_{u}^{2}\end{array}\right].$
(15) $\{{\bf l}_{itj}\mid{\bf x}_{i},U_{i}\}\sim\sum_{r\geq 1}\pi_{r}^{(tj)}({\bf x}_{i})\mathcal{N}(\gamma_{0r}+\gamma_{1r}U_{i},\sigma^{(l|u)2}_{r}),$
where the location and scale parameters are
$\gamma_{0r}+\gamma_{1r}U_{i}=\eta_{r}^{(tj)}+\frac{\sigma^{(lu)}}{\sigma_{u}^{2}}(U_{i}-\mu_{u}),$
$\sigma^{(l|u)2}_{r}=\frac{1}{\tau_{r}^{(tj)}}+\frac{\sigma^{(lu)}}{\sigma_{u}^{2}}.$
Property 2.
Assuming the unmeasured variable $U$ is normally distributed, the treatment-specific factors ${\bf L}_{t}$, for each $t$, are defined as in (3)-(5), and the assumptions of Eq.(11) hold, then each component of the mixture $r\geq 1$, has: (i) a finite expected value of the location parameter, that depend on the expected value of the unmeasured variable $U$, and (ii) the scale parameter has the scale parameter of the conditional distribution (15) as a lower bound. Specifically,
$(i)\quad{\mathds{E}}[\eta_{k}]=\gamma_{0r}+\gamma_{1r}\mu_{u}={\mathds{E}}[{\bf l}_{tj}\mid{\bf x},U],$
(16) $(ii)\quad\frac{1}{\tau_{r}^{(tj)}}=\sigma^{(l|u)2}_{r}+\frac{\sigma^{(lu)}}{\sigma_{u}^{2}}\geq\sigma^{(l|u)2}_{r}.$
Therefore, the probability distribution of $\{L_{tj}\mid{\bf X}\}$ has same mean as $\{L_{tj}\mid{\bf X},U\}$ and an equal or larger variance. Substituting this result into Eq. (1), we conclude that the casual effects—the SATE—are unbiased under the assumptions defined in Section Causal Estimands and Identifying Assumptions.
Scenario 1 in Figure 1 corresponds to the case where factor scores and the unmeasured variable are independent, i.e., $\sigma^{(lu)}=0$. In this case, the marginal and conditional variances are the same, indeed the equation (16) can be simplified in $1/\tau_{r}^{(tj)}=\sigma^{(l|u)2}_{r}$. Conversely, Scenarios 2 and 3 assume $\sigma^{(lu)}>0$.
@@
For the factor loading matrix, we adopt the well-known shrinkage prior introduced by Bhattacharya and Dunson (2011). For each treatment-specific factor loading $\lambda_{tjh}$, where treatment level $t\in\{0,1\}$, the factor $j=1,\ldots,\infty$ and the outcome variable $h\in\{1\ldots p\}$:
$\lambda_{tjh}\mid k_{tjh},\iota_{th}\sim N\left(0,\theta_{tjh}^{-1}\iota_{th}^{-1}\right),$
$\theta_{tjh}\sim\operatorname{Gamma}\left(\nu_{t}/2,\nu_{t}/2\right),$
$\iota_{th}=\prod_{l=1}^{h}\delta_{tl},$
$\delta_{t1}\sim\operatorname{Gamma}\left(a_{1_{t}},1\right),\;\delta_{tl}\sim\operatorname{Gamma}\left(a_{2_{t}},1\right)\forall l\geq 2.$
@@ 4. Simulation Study
We consider four simulation scenarios. The first three explore different relationships between the unmeasured variable $U$ and the confounders ${\bf X}$, corresponding to the graphical representations in Figure 1. The fourth scenario is designed to closely mimic the real dataset used in our application in Section 5. Environmental Application. Details of simulation-generating processes are provided in Table 1.
@@
Specifically, in the first three scenarios, the treatment-specific factor scores $\{{\bf L}_{0},{\bf L}_{1}\}$ depend on both the confounders ${\bf X}$ and the unmeasured variable $U$, while the potential outcomes ${\bf Y}$ depend on $U$ only through the factors. In Scenario 1, shown on the left of Figure 1, ${\bf X}$ and $U$ are independent (i.e., $\sigma^{lu}=0$ in Property 2). Scenario 2 assumes that $U$ depends on ${\bf X}$, and Scenario 3 explores the reverse relationship, where the confounders ${\bf X}$ depend on $U$. These dependencies are highlighted by the red dotted arrows in the second and third representations of Figure 1. Accordingly to Scenario 2, we consider four independent confounders ${\bf X}$ and model $U$ as dependent on ${\bf X}$. In contrast, Scenario 3 is modeled with an independent normal distribution for $U$; also the first two confounders in ${\bf X}$ are normally distributed with mean given by a linear function of $U$. Scenario 4 uses real-data information for the 28 confounders and the treatment assignment. Factor loadings are estimated with our model applied to real data. Factor scores and outcomes were simulated according to our proposed model.
For each of the four scenarios, we generated 50 collections of datasets. Causal effects are estimated using the R package bartCause to estimate with the BART model, the bcf code available on GitHub for BCF estimation, and our own Gibbs sampler for the factor models, available on GitHub at dafzorzetto/BayesCausalFactor. The same codebase includes the implementation of the standard factor model.
We evaluate the performance of the models according to the bias mean and square error (MSE) of SATEk  estimation for each $k=\{1,\dots,10\}$ in Scenarios 1–3, and $k=\{1,\dots,27\}$ in Scenario 4. The results are summarized in Figure 2. Our proposed model (CausalFA) outperforms the competitors in both bias and MSE, showing unbiased estimations of causal effects and a smaller variability. Specifically, BART and BCF exhibit higher values for the MSE, with particular ample boxplots for the BART. The factor model with standard normal prior (StandardFA) estimates causal effects with bias, highlighting a critical weakness of standard prior for factor scores and the benefit of the DDP prior.
Our proposed model is the only one among the four we considered to provide an unbiased estimation of the causal effects for all the elements in the multivariate outcomes across the simulated scenarios, demonstrating its ability to adjust the estimation for the unmeasured variable through the latent factors and their prior.
@@ 5. Environmental Application
Following Krasovich Southworth et al. (2025), who report strong temporal correlations in the data, we focus on a specific time period—specifically, the summer months—when wildfires are more frequent. Moreover, instead of using the prediction of the level of pm2.5 induced by wildfire smoke, our treatment variable is defined as the binary presence or absence of wildfire smoke.
@@
The data analyzed are merged from different sources, following a strategy similar to Krasovich Southworth et al. (2025). We consider the $280$ air pollution monitors across the United States, operated by the National Park Service and the U.S. Forest Service. The data are collected from the Environmental Protection Agency’s (EPA) pm2.5 Chemical Speciation Network (CSN) and the Interagency Monitoring of Protected Visual Environments (IMPROVE) program, and stored in the Federal Land Manager Environmental Database (F.E.D. IMPROVE, 2024). We select the same $27$ chemical species considered in Krasovich Southworth et al. (2025), divided in alkaline-earth metals (Magnesium, Calcium, and Strontium), alkali metals (Sodium, Rubidium, and Potassium), transition metals (Chromium, Nickel, Vanadium, Copper, Iron, Zinc, and Manganese), metalloids (Arsenic and Silicon), other metals (Lead, Aluminum, and Titanium), nonmetals (Selenium, Nitrate, Sulfate, Sulfur, and Phosphorus), halogens (Chlorine and Bromine), and Organics (elemental Carbon and organic Carbon).
@@
Wildfire smoke pm2.5 concentration are derived from the daily prediction of Childs et al. (2022), available from the GitHub repository echolab-stanford/daily-10km-smokePM. We define the treatment variable as $T=1$ for “smoke days” with non-zero predictions and $T=0$ for “non-smoke days” with zero predicted wildfire smoke pm2.5. The prediction resolution in the grid is $10Km$ (i.e., each unit aggregates the information over $10Km\times 10Km$ area), therefore we match the location of each air pollution monitor with the closest point of the grid.
@@
Study Design
Our analysis focuses on July–September 2014, when wildfire smoke is most prevalent in the United States (Krasovich Southworth et al., 2025). The observational unit is a monitor-day. Since chemical concentrations are measured every three days, the dataset contains $7,467$ units.
The multivariate outcome is the concentration of $27$ chemical species, measured in $\mu g/m^{3}$ and transformed on the natural logarithmic scale. Confounders include weather information, census data, type of monitor indication, month, latitude and longitude of monitor location, and type of monitor location (rural, city or intermediate).
Initially we have $5,754$ monitor day units assigned in the control group —-that is, no wildfire smoke day at that monitor location —-and $1,713$ in the treated group —that is, exposed to wildfire smoke. However, we use matching before running our model to make our analyses as robust as possible with respect to the potential measured confounding bias, using confounders to measure the similarity of the units in the two groups. This is common in observational studies (Rosenbaum and Rubin, 1983) and in research on air pollution effects on health (see, e.g., Lee et al., 2021; Zorzetto et al., 2024).
We employ a 1-to-1 nearest neighbor propensity score matching, yielding a balanced sample of $3,426$ units, and improving the balance of the covariates. The reduction in units is due to the different sample sizes of the treated and control groups in the original data, and 1-to-1 matching creates a sample with the same size for the treated and control groups. The causal effects, estimated in the following section, target the sample obtained through the matching procedure.
@@ Results
All four models consistently identify a positive causal effect of wildfire smoke on zinc, bromine, and the two organic elements—elemental and organic carbon—indicating that wildfire smoke increases the levels of these chemicals. This finding aligns with previous studies: Young and Jan (1977) and Odigie and Flegal (2014) observed increased zinc concentrations following wildfire smoke in California, Liu et al. (2014) reported elevated carbon levels and their negative impact on ecosystems, and Li et al. (2023) found higher bromine concentrations during wildfires in urban areas. However, the models do not fully agree on the detection of significant negative causal effects, i.e. a reduction in chemical concentrations due to wildfire smoke, particularly for strontium, sodium, nickel, titanium, silicon and aluminum. This discrepancy arises from the wide credible intervals produced by BART, which often include zero, and, in some cases, by BCF as well, limiting detection of significant effects. As shown in the simulation study, these two models tend to estimate wide variability, particularly in highly correlated outcomes. Similarly, for positive effects, only our proposed model has credible intervals that do not include zeros for cooper, lead, sulfate, and phosphorum. This is consistent with previous studies, as Young and Jan (1977) and Odigie and Flegal (2014) observed notable increases in these metals, while Spencer and Hauer (1991) reported elevated phosphorus concentrations following wildfires.
@@
The different distributions of the causal effects identified by the four models can be better understood by examining the factor loading matrix, reflecting the correlations among the 27 chemical species. Our Bayesian causal factor model framework enables us to analyze these relationships and assess how wildfire smoke alters them by estimating treatment-specific factors—i.e., distinct factors for the presence and absence of wildfire smoke. After applying a varimax transformation (Kaiser, 1960), we identify $3$ factors for each treatment level, explaining a total of $85\%$ and $88\%$ of the variance, respectively for wildfire smoke and non-smoke days (Figure 4). In both treatment levels, the first factor captures strong correlations among transition metals, metalloids, and other metals. Notably, titanium, silicon, and aluminum, chemicals whose concentrations drive the Factor 1, are detected to have significant wildfire smoke effects by only the factor models (CausalFA and StandardFA), highlighting their ability to include correlation across the outcome components and narrow uncertainty and more robust conclusions.
Factor 2 captures strong correlations among non-metals, while Factor 3 reflects associations among organic elements. In both cases, wildfire smoke is estimated to increase factor levels. This may be plausibly related to wildfires burning natural materials such as forests, releasing elemental and organic carbon into the air Liu et al. (2014). Likewise, non-metals, commonly found in soil, exhibit strong positive correlations with wildfire events (Spencer and Hauer, 1991).
Furthermore, the presence of wildfire smoke appears to enhance correlations among alkaline-earth metals, transition metals, and metalloids within Factor 1, particularly among chemicals that share the same negative causal effect. Similarly, in Factor 2, wildfire smoke strengthens correlations among non-metals, while in Factor 3, it disrupts the correlation between organic chemicals and other elements, reinforcing their internal associations. These structural shifts underscore the utility of our model in uncovering the latent structure of treatment.
@@ 6. Discussion
By combining these two contributions, the proposed Bayesian causal regression factor model demonstrates superior performance in causal effect estimation compared to state-of-the-art methods in both the factor analysis and causal inference domains, also correcting the bias in the causal effect estimates. Specifically, it outperforms standard factor models that rely on Gaussian priors for factor scores, which fail to characterize heterogeneity in factor scores and consequently impute the missing outcome. At the same time, it outperforms established methods in causal inference such as BART and BCF, which, despite their flexibility, lack the ability to share latent structure across correlated outcome components. Moreover, BART and BCF exhibit higher MSE values, likely due to their independent estimation of each causal effect. However, future extensions could enhance our causal factor model by leveraging the flexibility of BART priors Chipman et al. (2010). Additionally, alternative priors can be considered for both the covariate regression—such as mixtures or spike-and-slab priors George and McCulloch (1993)—and the factor loadings, including the Bayesian cumulative shrinkage prior by Legramanti et al. (2020) and the generalized sparse prior of Schiavon et al. (2022).
@@
While the proposed approach can be broadly applied to various real-world datasets for studying the effects of treatments or exposures on outcomes composed of correlated elements, our application in this paper focuses on a critical environmental issue. Specifically, we aim to estimate the causal effect of wildfire smoke on the chemical composition of pm2.5, motivated by the significant role that fine particulate exposure plays in human health. Our findings not only identify the chemicals that are positively or negatively affected by wildfire smoke but also reveal significant effects that other models fail to detect, due to its ability to borrow strength across outcomes. Moreover, our method effectively identifies the latent factors that drive the behavior of chemical components across different types of metals and non-metals, capturing treatment-specific patterns across classes of metals, nonmetals, and organics. This capability enriches our understanding of how wildfire smoke alters the composition of air quality in complex and systematic ways. Further research could extend this work by modeling multivariate outcomes under exposure to wildfire smoke over time, with methodologies similar to Samartsidis et al. (2020, 2024) or by accounting for the positive continuous nature of the prediction of pm2.5 induced by wildfire smoke, as discussed in Krasovich Southworth et al. (2025), though in a non-causal framework.
@@
We hope that this study highlights the importance of factor scores—not only in causal inference, as supported by recent studies such as Zorzetto et al. (2024) and Bortolato and Canale (2024)—and indicates the potential of factor models to quantify treatment effects in causal inference, particularly in settings where outcomes are high-dimensional and strongly correlated.

### v3

@@ 3. Bayesian Causal Factor Model
where $\Phi(\cdot)$ is the Probit function and $\{a_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})\}_{r\geq 1}$ have Gaussian distributions whose mean is a linear combination of the $p$ covariates ${\bf x}_{i}$.
@@
This prior distribution for treatment-specific factors (3) allows us to introduce a latent categorical variable $S_{it{\color[rgb]{0,0,0}h}}$, for each factor ${\color[rgb]{0,0,0}h}\in\{1,\dots,J_{t}\}$ treatment $t\in\{0,1\}$ unit $i\in\{1,\dots.n\}$, describing clusters of units defined by heterogeneous responses to the treatment level $t$. Assuming $\Pr(S_{it{\color[rgb]{0,0,0}h}}=r)=\pi_{r}^{(t{\color[rgb]{0,0,0}h})}({\bf x}_{i})$, then the prior distribution of the treatment-specific factors (3) can be rewritten, conditional on $S_{it{\color[rgb]{0,0,0}h}}$, as
(6) $\{{\bf l}_{it{\color[rgb]{0,0,0}h}}\mid S_{it{\color[rgb]{0,0,0}h}}=r,\boldsymbol{\omega}\}\sim\mathcal{N}(\eta_{r}^{(t{\color[rgb]{0,0,0}h})},1/\tau_{r}^{(t{\color[rgb]{0,0,0}h})}).$
For each treatment $t$ and cluster $r$, we assume $\tau_{r}{th}$ fixed to 1 and the prior distribution for the location parameter as $\eta_{r}^{(th)}\sim\mathcal{N}(0,1)$.
@@
The proposed prior for treatment-specific factor scores offers several strengths, as outlined in Section The Role of the DDP in the Factor Scores. Additionally, it serves as a powerful tool for inferring information about the relationships among the multiple outcomes, as governed by unobserved features $U$. A detailed discussion of these statistical guarantees is provided in the Supplementary Material.
@@
For the factor loading matrix, we adopt the well-known shrinkage prior introduced by Bhattacharya and Dunson (2011). For each treatment-specific factor loading $\lambda_{tjh}$, where treatment level $t\in\{0,1\}$, the factor $h=1,\ldots,\infty$ and the outcome variable $j\in\{1\ldots p\}$:
$\lambda_{tjh}\mid k_{tjh},\iota_{th}\sim N\left(0,\theta_{tjh}^{-1}\iota_{th}^{-1}\right),\;\theta_{tjh}\sim\operatorname{Gamma}\left(\nu_{t}/2,\nu_{t}/2\right),$
$\iota_{th}=\prod_{l=1}^{h}\delta_{tl},\;\delta_{t1}\sim\operatorname{Gamma}\left(a_{1_{t}},1\right),\;\delta_{tl}\sim\operatorname{Gamma}\left(a_{2_{t}},1\right)\forall l\geq 2.$
@@ 4. Simulation Study
We consider four simulation scenarios. The first three explore different relationships between the unmeasured variable $U$ and the confounders $\bf X$}, corresponding to the graphical representations in Figure 1. The fourth scenario is designed to closely mimic the real dataset used in our application in Section 5. Environmental Application. Details of simulation-generating processes are provided in Table 1 and fully described in the Supplementary Materials.
@@
For each of the four scenarios, we generated 100 collections of datasets. Causal effects are estimated using the R package bartCause to estimate with the BART model, the bcf code available on GitHub for BCF estimation, and our own Gibbs sampler for the factor models, available on GitHub at dafzorzetto/BayesCausalFactor. The same codebase includes the implementation of the standard factor model.
We evaluate the performance of the models according to the bias mean and square error (MSE) of SATEk  estimation for each $k=\{1,\dots,10\}$ in Scenarios 1–3, and $k=\{1,\dots,27\}$ in Scenario 4. The results are summarized in Figure 2. Our proposed model (CausalFA) outperforms the competitors in both bias and MSE, showing unbiased estimations of causal effects and a smaller variability. Specifically, BART and BCF exhibit higher values for the MSE, with particular ample boxplots for the BART. The factor model with standard normal prior (StandardFA) estimates causal effects with bias, highlighting a critical weakness of standard prior for factor scores and the benefit of the DDP prior. Only our model provides unbiased causal effect estimates across all simulated scenarios, effectively adjusting for unmeasured variables via latent factors and their priors. Moreover, Figure C.2 in the Supplementary Materials demonstrates good coverage properties for our proposed model.
Additional simulation setting are analyzed in the Supplementary Materials, Appendix C.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@ 5. Environmental Application
The data analyzed are merged from different sources, following a strategy similar to Krasovich Southworth et al. (2025). We consider the $280$ air pollution monitors across the United States, operated by the National Park Service and the U.S. Forest Service. The data are collected from the Environmental Protection Agency’s (EPA) pm2.5 Chemical Speciation Network (CSN) and the Interagency Monitoring of Protected Visual Environments (IMPROVE) program, and stored in the Federal Land Manager Environmental Database (F.E.D. IMPROVE, 2024). We select the same $27$ chemical species considered in Krasovich Southworth et al. (2025), divided in alkaline-earth metals, alkali metals, transition metals, metalloids, other metals, nonmetals, halogens, and Organics. The full list of chemical species is provided in the Supplementary Materials.
@@
Wildfire smoke pm2.5 concentration are derived from the daily prediction of Childs et al. (2022), available from the GitHub repository echolab-stanford/daily-10km-smokePM. We define the binary treatment variable “smoke day” as follows: for each monitor-day unit, treatment level $t=1$ (a “smoke day”) indicates that non-zero smoke-related pm2.5 is predicted at that monitor location for that day. Conversely $t=0$ (a “non-smoke day”) indicates that no smoke-related pm2.5 is predicted. Further details are in the Supplementary Materials.
In our analysis, SUTVA assumes that the wildfire smoke pm2.5 concentration at a given location is only dictated by the smoke at that location and not by smoke at other locations. Essentially, once we know the wildfire smoke at a given location, whichever wildfire(s) it might have originated from or the smoke at neighboring locations is not relevant to the chemicals at that location.
@@
Our analysis focuses on July–September 2014, when wildfire smoke is most prevalent in the United States (Krasovich Southworth et al., 2025). The observational unit is a monitor-day, for a total of $7,467$ units. The study design, described in detail in the Supplementary Materials, employs a 1-to-1 nearest neighbor propensity score matching, yielding a balanced sample of $3,426$ units, and improving the balance of the covariates. The causal effects, estimated in the following section, target the sample obtained through the matching procedure.
@@ Results
All four models consistently identify a positive causal effect of wildfire smoke on zinc, bromine, and the two organic elements—elemental and organic carbon—indicating that wildfire smoke increases the levels of these chemicals. This finding aligns with previous studies: Odigie and Flegal (2014) observed increased zinc concentrations following wildfire smoke in California, Liu et al. (2014) reported elevated carbon levels and their negative impact on ecosystems, and Li et al. (2023b) found higher bromine concentrations during wildfires in urban areas. However, the models do not fully agree on the detection of significant negative causal effects, i.e. a reduction in chemical concentrations due to wildfire smoke, particularly for strontium, sodium, nickel, titanium, silicon and aluminum. This discrepancy arises from the wide credible intervals produced by BART, which often include zero, and, in some cases, by BCF as well, limiting detection of significant effects. As shown in the simulation study, these two models tend to estimate wide variability, particularly in highly correlated outcomes. Similarly, for positive effects, only our proposed model has credible intervals that do not include zeros for cooper, lead, sulfate, and phosphorum. This is consistent with previous studies, as Odigie and Flegal (2014) observed notable increases in these metals, while Spencer and Hauer (1991) reported elevated phosphorus concentrations following wildfires.
@@
The different distributions of the causal effects identified by the four models can be better understood by examining the factor loading matrix, reflecting the correlations among the 27 chemical species. Our Bayesian causal factor model framework enables us to analyze these relationships and assess how wildfire smoke alters them by estimating treatment-specific factors—i.e., distinct factors for the presence and absence of wildfire smoke. After applying a varimax transformation (Kaiser, 1960), we identify $3$ factors for each treatment level, explaining a total of $85\%$ and $88\%$ of the variance, respectively for wildfire smoke and non-smoke days (Figure 4). In both treatment levels, the first factor captures strong correlations among transition metals, metalloids, and other metals. Notably, titanium, silicon, and aluminum, chemicals whose concentrations drive the Factor 1, are detected to have significant wildfire smoke effects by only the factor models (CausalFA and StandardFA), highlighting their ability to include correlation across the outcome components and narrow uncertainty and more robust conclusions. Factor 2 captures strong correlations among non-metals, while Factor 3 reflects associations among organic elements. In both cases, wildfire smoke is estimated to increase factor levels. This may be plausibly related to wildfires burning natural materials such as forests, releasing elemental and organic carbon into the air Liu et al. (2014). Likewise, non-metals, commonly found in soil, exhibit strong positive correlations with wildfire events (Spencer and Hauer, 1991). Furthermore, the presence of wildfire smoke appears to enhance correlations among alkaline-earth metals, transition metals, and metalloids within Factor 1, particularly among chemicals that share the same negative causal effect. Similarly, in Factor 2, wildfire smoke strengthens correlations among non-metals, while in Factor 3, it disrupts the correlation between organic chemicals and other elements, reinforcing their internal associations. These structural shifts underscore the utility of our model in uncovering the latent structure of treatment.
@@ 6. Discussion
By combining these two contributions, the proposed Bayesian causal regression factor model demonstrates superior performance in causal effect estimation compared to state-of-the-art methods in both the factor analysis and causal inference domains, also correcting the bias in the causal effect estimates. Specifically, it outperforms standard factor models that rely on Gaussian priors for factor scores, which fail to characterize heterogeneity in factor scores and consequently impute the missing outcome. At the same time, it outperforms established methods in causal inference such as BART and BCF, which, despite their flexibility, lack the ability to share latent structure across correlated outcome components. Moreover, BART and BCF exhibit higher MSE values, likely due to their independent estimation of each causal effect. However, future extensions could enhance our causal factor model by leveraging the flexibility of BART priors Chipman et al. (2010). Additionally, alternative priors can be considered for both the covariate regression—such as mixtures—and the factor loadings, including the generalized sparse prior of Schiavon et al. (2022).
@@
While the proposed approach can be broadly applied to various real-world datasets for studying the effects of treatments or exposures on outcomes composed of correlated elements, our application in this paper focuses on a critical environmental issue. Specifically, we aim to estimate the causal effect of wildfire smoke on the chemical composition of pm2.5, motivated by the significant role that fine particulate exposure plays in human health. Our findings not only identify the chemicals that are positively or negatively affected by wildfire smoke but also reveal significant effects that other models fail to detect, due to its ability to borrow strength across outcomes. Moreover, our method effectively identifies the latent factors that drive the behavior of chemical components across different types of metals and non-metals, capturing treatment-specific patterns across classes of metals, nonmetals, and organics. This capability enriches our understanding of how wildfire smoke alters the composition of air quality in complex and systematic ways. Further research could extend this work by modeling multivariate outcomes under exposure to wildfire smoke over time, with methodologies similar to Samartsidis et al. (2024) or by accounting for the positive continuous nature of the prediction of pm2.5 induced by wildfire smoke, as discussed in Krasovich Southworth et al. (2025), though in a non-causal framework.
@@
We hope that this study highlights the importance of factor scores—not only in causal inference, as recently shown by Zorzetto et al. (2024c) and Bortolato and Canale (2024)—and indicates the potential of factor models to quantify treatment effects in causal inference, particularly in settings where outcomes are high-dimensional and strongly correlated.